Ok i have to solve for x in this equation:
x^2 + (x+7)^2 = 13^2
this is my work but its wrong:
2x^2 + 49 = 169
2x^2 - 120 = 0
square root of (0- 4*-120*2) over 4
please help

Answer by edjones(8007)

x^2 + (x+7)^2 = 13^2
x^2+x^2+14x+49=169
2x^2+14x+49-169=0
2x^2+14x-120=0
2(x^2+7x-60)=0
(x+12)(x-5)=0
x=-12, x=5
.
Ed